OpenStudy (anonymous):

Which ordered pair in the form (x, y) is a solution of this equation? (x + 7)y = 10 A. (13, 2) B. (2, 1) C. (3, 1) D. (4, 1)

OpenStudy (anonymous):

i got b

OpenStudy (photon336):

let's see you mean x+7y = 10 ?

OpenStudy (photon336):

@Swagmadness251

OpenStudy (anonymous):

yeah the answer would be 2n1

OpenStudy (photon336):

should probably re-write this. x+7y = 10 7y = 10-x

OpenStudy (photon336):

then you can see what points satisfy the equation. by plugging in x, and if you get the same value for y then the point satisfies the equation.

OpenStudy (anonymous):

k

OpenStudy (photon336):

7(y) = 10-x let's try this one B (2,1) SO when x = 2 y must be 1 7(y) = 10-2 y = 8 so that point doesn't work. you must try again